OpenStudy (anonymous):

y+1=2/3(x+3) write in standard form. I don't know what to do with the +3 at the end.

OpenStudy (anonymous):

What is the question

OpenStudy (anonymous):

I have to write y+1=2/3(x+3) in standard form.

OpenStudy (anonymous):

y+2= -3(x-1) y+2= -3x+3 3x+y-1=0 so your final answer is: 3x+y-1

OpenStudy (anonymous):

Do you understand?

OpenStudy (amistre64):

almost a good example, just a few things off in it

OpenStudy (anonymous):

I don't think thats what teacher is looking for. Something like... -2x+3y=??? I don't know what to do with the +1 or the +3

OpenStudy (amistre64):

in your example: 3x+y-1=0 is good, or 3x+y=1 is fine

OpenStudy (amistre64):

standard form is Ax + By = C such that A is not negative.

OpenStudy (amistre64):

y+1=2/3(x+3) 3y+3=2(x+3) 3y+3=2x+6 -2x+3y=3 multiply thru by -1 so that we dont have a -2 starting out

OpenStudy (amistre64):

once we got here: 3y+3=2x+6 we could have just moved the y stuff over and the constants to the other side, since we have a positive x coefficient 3=2x-3y+6 -3=2x-3y